A construction company is considering whether to lease or buy equipment for its new 4-year project. If they buy the equipment please answer them all and mark the answers . thanks
Answer #1

Part1

Present worth (PW) of lease :

Cash flows are

Time cashflows

0 -125000

1 to 4 -210000 p.a. ( lease rental)

-210000× (1-.28)= $ -151200 p.a. after tax

4 125000

PW = -125000-151200×AF(17%,4) +125000× PVF(17%,4)

AF(17%,4) = annuity factor = (1-1.17^-4)/.17 = 2.74

PVF(17%,4)= present value factor= 1/1.17^4 = .5337

PW = -125000 - 151200× 2.74 +125000×.5337 = -$ 472576

PW of purchase

Cash flows

At time 0  

Initial investment = $630000

From time 1 to 4

After tax annual cost = 42000×(1-.28)= $30240

PV of this = 30240×AF(17%,4)= 30240×2.74= $82858

Depreciation as per MACRS Ads, Since question doesn't provide details which method to use we assume that SLM os being used as per this method annual Depreciation = 630000/4= $157500

Annual depreciation tax shield(DTS) = tax rate× 157500= .28×157500= $44100

PV of this Tax shield = 44100×2.74= $120834

At time 4

Sale of equipment = $378000

Capital gain= 378000 because ads SLM depreciate the value toll 0

Capital gains tax= 378000×15%= $56700

After tax sale= 378000- 56700= $321300

PV of this = 321300×PVF(17%,4)= 321300×.5337= 171478

Now the PW for purchase is given by

PW= PV of DTS+ PV of after tax sale - PV of after tax cost - intial cost

= 120834 +171478-82858-630000= - $420546

Since PW of total cost in less in option 2, we should go with this i.e. we should buy the equipment rather taking it on lease.